Getty Company expects sales for the first three months of next year to be $180,000, $245,000 and $275,000, respectively. Getty expects 40 percent of its sales to be cash and the remainder to be credit sales. The credit sales will be collected as follows: 10 percent in the month of the sale and 90 percent in the following month.

Compute a schedule of Getty’s cash receipts for the months of February and March.
Answer #1

Answer

--Schedule

Febraury March
Cash from Cash sales $98,000 $110,000
Cash from Credit sales of:
January $97,200
Febraury $14,700 $132,300
March $16,500
TOTAL $209,900 $258,800

--Working

Febraury March
Cash from Cash sales =245000*40% =275000*40%
Cash from Credit sales of:
January =180000*60%*90%
Febraury =245000*60%*10% =245000*60%*90%
March =275000*60%*10%